# SAMPO: Scale-wise Autoregression with Motion PrOmpt for generative world models

### v1

@@ 1 Introduction
Building a world model that can simulate the physical environment and respond to the actions of agents is a central challenge on the path to artificial general intelligence (AGI) [2, 11, 73, 34, 20, 55, 21, 22]. Recently, video generation has been integrated into world models, enabling models to generate future frames based on agent actions, simulating dynamic environments and making it possible for agents to anticipate outcomes and make informed decisions [25, 7, 50, 4]. Despite growing progress, designing a world model that is simultaneously high-fidelity, temporally consistent, and efficiently scalable remains an open problem.
@@ 2.2 Motion Prompt for Visual Dynamics Modeling
Visual prompt have emerged as a lightweight alternative to architectural changes for guiding multimodal models [44, 62, 61, 65, 59]. Early efforts introduce coarse overlays or fine‑grained masks to input images, steering large vision–language models toward target regions [44, 62, 65]. While effective for object localization [61, 59], these prompts encode only static spatial cues and fail to capture object motion, limiting their suitability for world model learning and control. To address this limitation, recent works have introduced motion-aware prompting techniques that explicitly encode spatio-temporal dynamics [16, 70, 35, 57]. Motion prompting [16] controls video diffusion models using sparse or dense motion tracks, enabling realistic and controllable object and camera dynamics. TraceVLA [70] overlays tracked trajectories [32] as visual prompts to inject temporal context into vision–language action models without architectural changes. Complementarily, MoVideo [35] integrates optical flow and depth features to enhance motion fidelity and temporal consistency in video generation. These methods demonstrate that motion-aware prompt can enhance dynamic fidelity without compromising efficiency. However, they often target generative tasks or depend on offline trajectories. In contrast, we propose an online motion prompting scheme that integrates with interactive world models for efficient and physically constrained visual control.
@@ 4.3 Ablation Studies & Model Analysis
Table 5: Performance and speed trade-off. We benchmark FVD, PSNR and inference speed using one A800 GPU with a batch size of 16 on the BAIR.
Method Spatial Scales FVD↓ PSNR↑ Inference Time↓
AR138M [58] - 60.8 24.5 9.05 s / vid.
SAMPO-S [1, 2, 4, 8] 80.7 23.1 1.61 s / vid
SAMPO-S [1, 2, 4, 8, 10] 73.1 24.3 3.27 s / vid
SAMPO-S [1, 2, 3, 4, 5, 6] 55.5 26.7 2.06 s / vid.
@@ A.1 Details of Network
Transformer. The SAMPO transformer utilizes a scalable decoder-only architecture, inspired by GPT-2, to efficiently model spatiotemporal dynamics. Unlike iVideoGPT [58], which relies on specialized tokens for frame segmentation, we initialize each frame with a single start token [S]. This token serves a dual function: it marks the beginning of intra-frame autoregressive generation and naturally defines the temporal boundaries between consecutive frames. Fixed 1D sine-cosine positional embeddings are applied to encode temporal dynamics, following standard practices in vision Transformers [40, 12, 26]. We design a set of models with different sizes, as illustrated in Tab. B. Model scaling adheres to a linear relationship between depth $d$, width $w$, head count $h$, and dropout rate $dr$:

### v2

@@ 1 Introduction
Building a world model that can simulate the physical environment and respond to the actions of agents is a central challenge on the path to artificial general intelligence (AGI) [2, 11, 77, 33, 52, 72, 73]. Recently, video generation has been integrated into world models, enabling models to generate future frames based on agent actions, simulating dynamic environments and making it possible for agents to anticipate outcomes and make informed decisions [24, 7, 50, 4]. Despite growing progress, designing a world model that is simultaneously high-fidelity, temporally consistent, and efficiently scalable remains an open problem.
@@ 2.2 Motion Prompt for Visual Dynamics Modeling
Visual prompt have emerged as a lightweight alternative to architectural changes for guiding multimodal models [43, 64, 63, 67, 61]. Early efforts introduce coarse overlays or fine‑grained masks to input images, steering large vision–language models toward target regions [43, 64, 67, 53]. While effective for object localization [63, 61], these prompts encode only static spatial cues and fail to capture object motion, limiting their suitability for world model learning and control. To address this limitation, recent works have introduced motion-aware prompting techniques that explicitly encode spatio-temporal dynamics [16, 74, 34, 59]. Motion prompting [16] controls video diffusion models using sparse or dense motion tracks, enabling realistic and controllable object and camera dynamics. TraceVLA [74] overlays tracked trajectories [31] as visual prompts to inject temporal context into vision–language action models without architectural changes. Complementarily, MoVideo [34] integrates optical flow and depth features to enhance motion fidelity and temporal consistency in video generation. These methods demonstrate that motion-aware prompt can enhance dynamic fidelity without compromising efficiency. However, they often target generative tasks or depend on offline trajectories. In contrast, we propose an online motion prompting scheme that integrates with interactive world models for efficient and physically constrained visual control.
@@ 4.2 Planning and Reinforcement Learning
Table 5: Performance and speed trade-off. We benchmark FVD, PSNR, average success rate on VP2 and inference speed using one A800 GPU with a batch size of 16 on the BAIR.
Method Spatial Scales FVD↓ PSNR↑ Avg. Success↑ Inference Time↓
AR138M [60] - 60.8 24.5 70.1 9.05 s / vid.
SAMPO-S [1, 2, 4, 8] 80.7 23.1 72.2 1.61 s / vid
SAMPO-S [1, 2, 4, 8, 10] 73.1 24.3 71.3 3.27 s / vid
SAMPO-S [1, 2, 3, 4, 5, 6] 55.5 26.7 70.6 2.06 s / vid.
4.3 Ablation Studies & Model Analysis
@@ 5 Conclusion and Discussion
Acknowledgement
This work was supported in part by the National Key Research and Development Project under Grant 2024YFB4708100, National Natural Science Foundation of China under Grants 62088102, U24A20325 and 12326608, and Key Research and Development Plan of Shaanxi Province under Grant 2024PT-ZCK-80.
@@ A.1 Details of Network
Transformer. The SAMPO transformer utilizes a scalable decoder-only architecture, inspired by GPT-2, to efficiently model spatiotemporal dynamics. Unlike iVideoGPT [60], which relies on specialized tokens for frame segmentation, we initialize each frame with a single start token [S]. This token serves a dual function: it marks the beginning of intra-frame autoregressive generation and naturally defines the temporal boundaries between consecutive frames. Fixed 1D sine-cosine positional embeddings are applied to encode temporal dynamics, following standard practices in vision Transformers [39, 12, 25, 44]. We design a set of models with different sizes, as illustrated in Tab. B. Model scaling adheres to a linear relationship between depth $d$, width $w$, head count $h$, and dropout rate $dr$: